# The Yōkai Learning Environment: Tracking Beliefs Over Space and Time

### v1

@@ C.2.2 GCN, GATv2 and the pre-embedding of Nodes
We use two graph neural networks in our work, one based on graph convolutions Kipf and Welling [2016] and one based on the attention mechanism Brody et al. [2021]. We pick these two as they stem from two different families of graph learning architectures and they thus are representative of a larger body of work. In our work, both methods observe the graph directly as described in the paper. There are two additional implementation details to note, though. First, both graph architectures observe an input graph that features a virtual node, similar to the hub-nodes of Abdessaied et al. [2024] or the [VNode] token of Ying et al. [2021]. A virtual node is an empty node that connects to all other nodes in the graph. It solely serves the purpose of reducing hop distances between nodes. The motivation for employing them in the YLE is that since cards can be distant from each other in terms of hop distance, information from one card might not propagate to the next during message passing. I.e. imagine the GNN observes two blue cards on opposite sides of the current card cluster. If the number of message-passing rounds is too low, this information might never be aggregated during message-passing. In RL, neural networks tend to be comparatively small, which is also true in our work. This also limits the number of message-passing steps to be small (2 in our case). With a virtual node, however, this is not an issue since the virtual node aggregates information from all other nodes in one single message-passing step. Second, our work uses pre-embeddings of node-feature modalities as described in Bortoletto et al. [2024d]. We take the node features of size $n\times f$ and embed them per modality via a separate dense encoder before then concatenating all of the resulting vectors into new node features shaped $n\times f^{\prime}$. $f^{\prime}$ depends on the size of these dense layers. We sketch the overall approach in Figure 13. In general, we found that embedding the node features per modality before using them in the GNN improved performance following earlier work, i.e. see Bortoletto et al. [2024d]. Specifically, for each modality $m\in M$ we train a dense layer $f_{m}$ that embeds the modality per node into an embedding $e_{m}\in\mathbb{R}^{b\times t\times n\times 2}$ for the current player embedding or $e_{m}\in\mathbb{R}^{b\times t\times n\times 4}$ for all other modalities where $b$ is the batch size, $t$ the time and $n$ the number of nodes. All embeddings are then concatenated into a node embedding $e_{\text{node}}=e_{\text{colour}}\oplus e_{\text{position}}\oplus e_{\text{locked}}\oplus e_{\text{seen}}\oplus e_{\text{id}}\oplus e_{\text{is\_current\_player}}\oplus e_{\text{substep}}$ with $e_{\text{node}}\in\mathbb{R}^{b\times t\times n\times f^{\prime}}$ with $f^{\prime}=26$ using the configuration described. We then add self-connections, the virtual node using a zero-vector $e_{\text{virt}}\in\{0\}^{f^{\prime}}$ and connect the virtual node to all real nodes before passing them to our graph encoder $f_{\text{GNN}}$. The output dimension of the graph encoder is $64$. For GATv2, we use $4$ heads. We apply $4$ graph layers in the relational encoder comparison. The resulting embedding is then concatenated along the node dimension and fed to the rest of the policy.
C.2.3 Relation Module
@@
Our work adapts the relation module of Zambaldi et al. [2018]. We only use the self-attention mechanism of their proposed relation module and skip the CNN encoder they propose alongside the self-attention layers. This is because our setting deals with an environment where agents observe the position of cards directly as opposed to the setting of Zambaldi et al. [2018] that deals with image inputs directly. Thus, in our relation module, non-local interactions between all cards are simply computed using self-attention directly. Similarly to the GNN encoders, we also pre-embed the modalities of all nodes using the same mechanism as discussed above. The resulting embedding is then also passed to the remaining policy. Our relational module is also applied $4$ times, features $4$ heads, and its output dimension per layer is $64$.
C.2.4 CNN
@@
Our work compares a CNN encoder and determines it to be the best-performing model. As described earlier, the YLE returns an image-like observation to be processed by our CNN. The exact size of the input tensor to the CNN depends on the YLE settings, but in 2-player 3x3 YLE, the input dimension is $9\times 10\times 13$, which we will use as a running example. It includes the same features as discussed above, with some minor changes. The first three channels include a one-hot encoding of cards’ colours, the next three an encoding of hint colours, followed by two channels for the position, one for whether the card is locked, one for whether the observing agent is the current player, one for the number of the current substep (1-4) and one that identifies the card via an identifier number. Our convolutional stack uses $4$ 2d convolutions with $64$ filters each and ReLU applied between them. The kernel size is $(3,3)$, we apply a stride $1$ and valid padding. The resulting embedding is flattened to produce the final embedding to be passed on to later layers.
C.2.5 LSTM and GRU
Next to the encoder architectures, our work compares several memory architectures. Here, memory architecture is used as a term to denote classical RNNs as well as structured state space or transformer Vaswani et al. [2017] architectures. LSTMs Hochreiter and Schmidhuber [1997] and GRUs Cho et al. [2014] are commonly used in RL and such are natural first approaches for our environment. They provide an implicit memory through their hidden state that is computed at every step from the current observation encoding and the previous hidden state. In both cases, we take the embedding generated by the encoder, $e_{\text{encoder}}$, pass it through a dense layer $f_{\text{pre\_rnn}}$ before finally applying either the LSTM or GRU on the resulting embedding. $f_{\text{pre\_rnn}}$ embeds $e_{\text{encoder}}$ to the hidden dimension of the RNN. As the hidden dimension of the RNN we pick $128$ in our work. We also experimented with using $256$ but found no difference in performance. We use orthogonal initialisation for both the GRU and the LSTM.
C.2.6 S5
As already described, we use S5 Smith et al. [2023] layers because of the great performance of structured state-space models on memory tasks in deep model-based RL Samsami et al. [2024]. Similar to LSTM and GRU, S5 only provides an implicit memory mechanism. We use a similar setup as with both the LSTM and GRU layers, i.e. we also use a $f_{\text{pre\_s5}}$ embedding layer that embeds $e_{\text{encoder}}$ to the hidden dimension of our S5 layer. We use $2$ S5 blocks, $1$ layer, place the LayerNorm Ba et al. [2016] before the S5 layer and use the activation function proposed in the original work Smith et al. [2023]. As before we use $128$ as the hidden dimension of S5.
C.2.7 TransformerXL
@@
Finally, as already introduced TransformerXL Dai et al. [2019] is an explicit memory mechanism that shows great performance in memory-intensive RL tasks when used correctly Parisotto et al. [2020]. Our usage strictly follows the one described by Parisotto et al. [2020] as implemented by Hamon [2024]. Similar to the other memory architectures, we also first pass the embedding received by the encoder through a dense layer $f_{\text{pre\_trxl}}$ before handing the resulting embedding to the transformer. While LSTM, GRU and S5 use $128$ as their hidden dimension, we only use a hidden dimension of $64$ for both the dense layers within the transformer as well as the projection layers in multi-head attention. This is to keep the number of parameters consistent between the architectures. We furthermore configure the transformer to have $8$ heads and to use gating with a gating bias of $2$. We use a single transformer layer. Finally, we set the length memory window to $32$. $32$ is long enough to recall all of the previous moves in the 3x3 YLE. Memories from a past game – as marked by the done signal – are masked. We apply standard positional encoding based on sinus waves.
C.3 ToMTrack
Agent Memory R $\uparrow$ SR in % $\uparrow$ NC $\uparrow$ SEE in % $\uparrow$
ToMTrack w/o $f_{\text{belief}}$ explicit $0.9\pm 0.0$.5) $31.2\pm 1.4$.8) $1.8\pm 0.0$.4) $0.0\pm 0.0$.0)
ToMTrack w/o action explicit $1.4\pm 0.0$.4) $41.2\pm 1.1$.7) $2.0\pm 0.0$.3) $0.0\pm 0.0$.0)
ToMTrack explicit $1.6\pm 0.0$.3) $48.3\pm 1.0$.7) $2.2\pm 0.0$.2) $0.0\pm 0.0$.0)
Table 7: ToMTrack ablations. We evaluate ToMTrack without the additional belief representation learning and the additional action encoding and find that only both together improve performance.
ToMTrack is a direct extension of the CNN-TrXL baselines and shares all the parameters mentioned earlier with both the CNN as well as the TransformerXL work. There are two extensions we add to arrive at ToMTrack. The first one – the auxiliary belief prediction task – is already mostly described in the main text. To generate the distribution over possible card colours, we add a colour prediction head $f_{\text{belief}}$ to our policy. $f_{\text{belief}}$ takes as input the embedding generated by the stack of dense layers and produces an output $e_{\text{belief}}\in\mathbb{R}^{b\times t\times|Y|c}$ that resembles the prediction over all cards and colours. For the 3x3 version $|Y|c=9\cdot 3=27$. To obtain the final predictions, we reshape $e_{\text{belief}}$ to be of shape $b\times t\times|Y|\times c$ and apply the Softmax function of the last dimension to get a prediction per card. The method only has one hyperparameter $\lambda$ – which balances the losses - and we pick $\lambda=0.1$ in this work. The second extension – the active action decoding – is realised via an additional input stream. Before passing the previous action to this stream, we first decode it into a reduced dimension that serves as a dense representation of that action of shape $4+|Y|+2+|H|+1$. The first $4$ bits encode the move type alias the substep. The next $|Y|$ bits encode the Yōkai card affected by the current action, if no card is affected, all are set to 0. The next $2$ bits encode the position affected by the action. For instance, if the previous active action was a move action, the resulting coordinates are encoded there. The next $|H|$ bits encode whether and if so which hint card is used in an action, and finally, the last bit encodes whether the previous action was to terminate the game. For a 2-player 3x3 YLE game $4+|Y|+2+|H|+1=20$. Note that this is a much denser representation compared to just encoding the raw number. The resulting encoding $e_{\text{action}}$ is processed by its own dense layer and then concatenated with the output of the encoder before being passed to the rest of the policy network. Note that we have ignored any hidden or memory state in the above formulation for notational convenience.
C.4 Other-Play in Yōkai
@@
The core idea of Other-Play (OP) is that during self-play training, agents learn to break symmetries of the environment in arbitrary ways that do not generalise to novel partners. In Hanabi, for instance, agents might form conventions around certain colours. OP presents agents with a special asymmetric domain randomisation that makes it impossible for agents to coordinate on breaking symmetries in certain ways. In Hanabi, this is achieved by a recolouring operation that shows each agent a different recolouring of the game state Hu et al. [2020].
@@
Note that these are not the only ways for agents to learn to break symmetries. Recent work Treutlein et al. [2021] showed that agents also might coordinate to break symmetries in arbitrary ways over multiple timesteps. It is impossible to list all possible classes of symmetries for an environment with enough complexity. This is a drawback of the OP algorithm as it requires them as input. In our work, we thus focus on showing that while Hanabi only requires breaking one symmetry for efficient ZSC while the YLE – due to its additional complexities – requires additional effort to improve ZSC performance, which is still far from optimal.
Appendix D Additional Probing Details
Over the duration of $1000$ games, we collect a probing dataset $\mathcal{D}=\{x^{(i)},y^{(i)}\}$. The probing datasets map the activations of an agent to a label $z$. A probe then is a function $g:f_{l}(x)\to\hat{z}$ and trained on $\mathcal{D}$ where $f$ is a neural network and $f_{l}(x)$ is the intermediate representation of $x$ at layer $l$. In our case, we train simple logistic regressions. We always probe before the action head of our agent. As highlighted in the main text, during all experiments, we train one and evaluate it on a hold-out test set $\mathcal{D}_{\text{test}}\subset\mathcal{D}$. Additionally, note that many of these tasks feature imbalanced datasets since the does-not-know label is far more likely to occur than any other label. We correct for that by weighing the loss per class and subsampling the does-not-know label to the largest other class. During all experiments, we train one probe per card – 9 in total – and average the scores computed on a hold-out test over all 9.
Appendix E Additional Experimental Results
E.1 Additional Training Curves
Figure 15: Return, success rate and numbers of clusters over time per encoder while given access to perfect memory.
Figure 16: Return, success rate and numbers of clusters over time per memory architecture.
Figure 17: Return, success rate and numbers of clusters over time for ToMTrack, ToMTrack-SAD, CNN-TrXL and the CNN-M architecture.
For completeness, we show how the most important metrics of our agents developed over time. Figure 15 shows the curves of the spatial-reasoning experiment, Figure 16 of the memory-reasoning experiment and lastly Figure 17 reports the same for ToMTrack.
E.2 Additional Zero-Shot Coordination Metrics
Figure 18: Zero-shot coordination performance measured by success-rate, average return and number of cards clustered. SP and OP+C are consistently outperformed by OP+C+R, showing that beyond cards’ colour, agents need to focus on other reasoning types.
Figure 19: We show training curve metrics for our agents trained with other-play to demonstrate convergence. Recall that these metrics measure self-play performance rather than ZSC. ZSC metrics are shown in Figure 6 and Figure 18.
In Figure 18 we show additional metrics for the ZSC evaluation in our main body of work. Training curves for our other-play agents are shown in Figure 19 to show convergence behaviour.
E.3 2- versus 4-player Games
Figure 20: ToMTrack in the two- and four-player 3x3 YLE. In four-player YLE more hints are available to be placed, and as such, we observe higher total rewards collected by the ToMTrack-4P agent, even though it wins fewer games and clusters fewer colours.
In Figure 20 we show how two- and four-player settings compare for our best model, ToMTrack.
@@ Appendix F Human Performance in Yōkai
We collected human game performance in a setting that adheres to the same setting that artificial agents face, i.e. participants played with 9 cards on a 9x9 grid in a two-player setting. Humans played in person. A pair of subjects was instructed on the rules verbally. Each pair of subjects was given a single training round to make sure that all participants were familiar with the rules. To make results as comparable as possible, humans were instructed to avoid any form of non-verbal communication. The study was conducted under an umbrella agreement of the relevant institutional review board. Compensation was also handled under the rules of the relevant institution.
@@ Appendix G Example Games
Figure 21: We show an example game in which ToMTrack wins in self-play. For each turn, we display the ground truth information at the top and the current observable information – essentially the agents’ observation – at the bottom. We also show how many card clusters exist at this point per turn. Hint cards are highlighted by a light-blue border. In turn 7, the agent finishes the final cluster of cards by correctly inferring that the card under the rbrb-hint must be red and not blue. In turn 8, agent 1 should have finished early for more reward, but did not recognise that fact and instead used the last hint card, thus finishing the game. Note that the agents tend to play quite conservatively. In turns 1 and 3, for instance, it gains knowledge of two separate green cards but only groups them in turn 5 after agent 1 places a corresponding gbgb-hint in turn 4.
@@
Figure 22: We show an example game in which ToMTrack loses in self-play. For each turn, we display the ground truth information at the top and the current observable information – essentially the agents’ observation – at the bottom. We also show how many card clusters exist at this point per turn. Hint cards are highlighted by a light-blue border. Note that agent 0 in turn 1 finds a blue group early and manages to form a cluster quickly. Agent 0 confirms this by turn 3. In turn 6, the game is still winnable. However, because both agents over-focused on establishing a common ground around the blue cluster, they fail to realise this and in turn 7, agent 0 then blunders and misplaces the green card. From there, the game is lost under all circumstances. Additionally, note that in turns 4 and 6, agent 1 establishes the identity of all red cards but fails to recall that in turn 6, therefore also misplaying.
@@
We show an example game in which ToMTrack wins in self-play in Figure 21 and one where it loses in Figure 22.
Appendix H Details on Other Attempts
While ToMTrack provides a mechanism for learning representations that represent card colours, it does not necessarily encourage the agents to maximise their shared or individual knowledge. Consequently, we were inspired to also experiment with a variant of ToMTrack that is not trained via an auxiliary task but rather uses $\mathcal{L_{\text{CE}}}$ as intrinsic motivation for both agents; ToMTrack-I. ToMTrack-I agents are encouraged to form an implicit belief model that aligns well with the true state of the environment by adding an intrinsic reward $r_{i}$ to the reward based on the negative cross-entropy loss of the actor classification head:
$r=r+\lambda_{\text{s}}*r_{\text{s}}+r_{\text{i}}\text{, where }r_{\text{i}}=-\mathcal{L}^{\text{CE}}_{a}.$ (9)
In this formulation, the agents thereby receive a higher reward when the belief model’s prediction loss $\mathcal{L}^{\text{CE}}_{a}$ is low. This approach is loosely inspired by Oguntola et al. [2023]. We originally assumed that this would achieve superior results but found that the intrinsic reward confused the agents to a point where they barely learned a useful policy altogether, falling below baseline performance (i.e. CNN-TrXL). Future work should study more sophisticated ideas based on this concept.
Agent Memory R $\uparrow$ SR in % $\uparrow$ NC $\uparrow$ SEE in % $\uparrow$
ToMTrack (SAD) explicit $0.7\pm 0.0$.4) $21.5\pm 1.3$.4) $1.5\pm 0.0$.4) $0.0\pm 0.0$.0)
Table 8: Additional ToMTrack-SAD results. This approach underperforms our baselines from Table 1 considerably and was thus excluded from the main paper.
Previous work on Hanabi has proposed the simplified action decoder (SAD) Hu and Foerster [2020]. SAD agents can take two actions during training that other agents can observe: a normal environment action and a greedy action. Other agents are allowed to observe both. The intuition is that during exploration, the true intentions of agents are blurred with explorative actions and thus less informative for others. Hu and Foerster [2020] implements their idea on top of epsilon-greedy Q-learning. Since during evaluation $\epsilon=0$, all actions are greedy and are given as input to both action inputs heads of other agents. During evaluation, $\epsilon=0$ and both actions are greedy. SAD showed improved performance in Hanabi, including the Bayesian action decoder Foerster et al. [2019]. We have adapted this idea as ToMTrack-SAD to our policy learning approach. Since we do not have an $\epsilon$ to set, we adopt their idea by also allowing agents to take two actions during training: One sampled from the agents’ produced action distribution is used as the environment action that gets executed, and one greedy action is determined as the legal action with the highest associated logit. The true environment action is duplicated during evaluation with similar reasoning.
ToMTrack-SAD is our extension of ToMTrack. In ToMTrack-SAD both actions are encoded as described earlier and concatenated before being fed to the same dense layer. Since we use MAPPO and no Q-learning approach as in the original work Hu and Foerster [2020], our agents pick their normal environment action given their current set of parameters $\theta_{a}$ in state $s_{t}$ via
$a_{\text{env}}\sim\pi(\cdot|s_{t},\theta_{a})$ (10)
and their second greedy action as
$a_{\text{greedy}}=\arg\max_{i}\pi(a_{i}|s_{t},\theta_{a}).$ (11)
Both are encoded and processed as described above when discussing ToMTrack. Note that we have ignored any hidden or memory state in the above formulation for notational convenience.
Results in Table 8 show that ToMTrack-SAD performs worse than CNN-TrXL or other baselines. We speculate that including greedy actions as input is uninformative in an environment with many actions like the YLE.

### v2

@@ D.2 Neural Network Architectures
We use two graph neural networks in our work, one based on graph convolutions Kipf & Welling (2016) and one based on the attention mechanism Brody et al. (2021). We pick these two as they stem from two different families of graph learning architectures and they thus are representative of a larger body of work. In our work, both methods observe the graph directly as described in the paper. There are two additional implementation details to note, though. First, both graph architectures observe an input graph that features a virtual node, similar to the hub-nodes of Abdessaied et al. (2024) or the [VNode] token of Ying et al. (2021). A virtual node is an empty node that connects to all other nodes in the graph. It solely serves the purpose of reducing hop distances between nodes. The motivation for employing them in the YLE is that since cards can be distant from each other in terms of hop distance, information from one card might not propagate to the next during message passing. I.e. imagine the GNN observes two blue cards on opposite sides of the current card cluster. If the number of message-passing rounds is too low, this information might never be aggregated during message-passing. In RL, neural networks tend to be comparatively small, which is also true in our work. This also limits the number of message-passing steps to be small (2 in our case). With a virtual node, however, this is not an issue since the virtual node aggregates information from all other nodes in one single message-passing step. Second, our work uses pre-embeddings of node-feature modalities as described in Bortoletto et al. (2024c). We take the node features of size $n\times f$ and embed them per modality via a separate dense encoder before then concatenating all of the resulting vectors into new node features shaped $n\times f^{\prime}$. $f^{\prime}$ depends on the size of these dense layers. We sketch the overall approach in Figure 9. In general, we found that embedding the node features per modality before using them in the GNN improved performance following earlier work, i.e. see Bortoletto et al. (2024c). Specifically, for each modality $m\in M$ we train a dense layer $f_{m}$ that embeds the modality per node into an embedding $e_{m}\in\mathbb{R}^{b\times t\times n\times 2}$ for the current player embedding or $e_{m}\in\mathbb{R}^{b\times t\times n\times 4}$ for all other modalities where $b$ is the batch size, $t$ the time and $n$ the number of nodes. All embeddings are then concatenated into a node embedding $e_{\text{node}}=e_{\text{colour}}\oplus e_{\text{position}}\oplus e_{\text{locked}}\oplus e_{\text{seen}}\oplus e_{\text{id}}\oplus e_{\text{is\_current\_player}}\oplus e_{\text{substep}}$ with $e_{\text{node}}\in\mathbb{R}^{b\times t\times n\times f^{\prime}}$ with $f^{\prime}=26$ using the configuration described. We then add self-connections, the virtual node using a zero-vector $e_{\text{virt}}\in\{0\}^{f^{\prime}}$ and connect the virtual node to all real nodes before passing them to our graph encoder $f_{\text{GNN}}$. The output dimension of the graph encoder is $64$. For GATv2, we use $4$ heads. We apply $4$ graph layers in the relational encoder comparison. The resulting embedding is then concatenated along the node dimension and fed to the rest of the policy.
Relation Module
@@
Our work adapts the relation module of Zambaldi et al. (2018). We only use the self-attention mechanism of their proposed relation module and skip the CNN encoder they propose alongside the self-attention layers. This is because our setting deals with an environment where agents observe the position of cards directly as opposed to the setting of Zambaldi et al. (2018) that deals with image inputs directly. Thus, in our relation module, non-local interactions between all cards are simply computed using self-attention directly. Similarly to the GNN encoders, we also pre-embed the modalities of all nodes using the same mechanism as discussed above. The resulting embedding is then also passed to the remaining policy. Our relational module is also applied $4$ times, features $4$ heads, and its output dimension per layer is $64$.
CNN
@@
Our work compares a CNN encoder and determines it to be the best-performing model. As described earlier, the YLE returns an image-like observation to be processed by our CNN. The exact size of the input tensor to the CNN depends on the YLE settings, but in 2-player 9C YLE, the input dimension is $9\times 10\times 13$, which we will use as a running example. It includes the same features as discussed above, with some minor changes. The first three channels include a one-hot encoding of cards’ colours, the next three an encoding of hint colours, followed by two channels for the position, one for whether the card is locked, one for whether the observing agent is the current player, one for the number of the current substep (1-4) and one that identifies the card via an identifier number. Our convolutional stack uses $4$ 2d convolutions with $64$ filters each and ReLU applied between them. The kernel size is $(3,3)$, we apply a stride 11 and valid padding. The resulting embedding is flattened to produce the final embedding to be passed on to later layers.
TransformerXL
@@
Next to the encoder architectures, our work compares memory architectures. TransformerXL (Dai et al., 2019) is an explicit memory mechanism that shows great performance in memory-intensive RL tasks (Parisotto et al., 2020). Our usage strictly follows the one described by Parisotto et al. (2020) as implemented by Hamon (2024). Similar to our GRU architecture, we first pass the embedding received by the encoder through a dense layer $f_{\text{pre\_trxl}}$ before handing the resulting embedding to the transformer. We use a hidden dimension of $64$ for both the dense layers within the transformer as well as the projection layers in multi-head attention. This is to keep the number of parameters consistent between the architectures. We furthermore configure the transformer to have $8$ heads and to use gating with a gating bias of $2$. We use a single transformer layer. Finally, we set the length memory window to $32$. $32$ is long enough to recall all of the previous moves in the 9C YLE. Memories from a past game are masked. We apply standard positional encoding based on sinus waves.
D.3 Additional Details for Other-Play in Yōkai
@@
The core idea of Other-Play (OP) is that during self-play training, agents learn to break symmetries of the environment in arbitrary ways that do not generalise to novel partners. In Hanabi, for instance, agents might form conventions around certain colours. OP presents agents with a special asymmetric domain randomisation that makes it impossible for agents to coordinate on breaking symmetries in certain ways. In Hanabi, this is achieved by a recolouring operation that shows each agent a different recolouring of the game state (Hu et al., 2020).
@@
Note that these are not the only ways for agents to learn to break symmetries. Recent work (Treutlein et al., 2021) showed that agents also might coordinate to break symmetries in arbitrary ways over multiple timesteps. It is impossible to list all possible classes of symmetries for an environment with enough complexity. This is a drawback of the OP algorithm as it requires them as input. In our work, we thus focus on showing that while Hanabi only requires breaking one symmetry for efficient ZSC while the YLE – due to its additional complexities – requires additional effort to improve ZSC performance, which is still far from optimal.
Appendix E Additional Probing Details
Figure 11: Probing results where only known card colours from openend cards where probed. We can see that agents start of representing card colours in a linearly decodeable way but as the game progresses and more cards are opened, accuracy decreases.
Figure 12: Probing results where only unknown card colours from unopenend cards where probed. As the game progresses card colours become increasingly linearly decodable but remain under a perfect baseline. Only OBL and OP represent unknown cards in a linear-decodable fashion above random chance.
Over the duration of $5,000$ games, we collect a probing dataset $\mathcal{D}=\{x^{(i)},y^{(i)}\}$. The probing datasets map the activations of an agent to a label $z$. A probe then is a function $g:f_{l}(x)\to\hat{z}$ and trained on $\mathcal{D}$ where $f$ is a neural network and $f_{l}(x)$ is the intermediate representation of $x$ at layer $l$. We always probe before the action head of our agent. As highlighted in the main text, during all experiments, we train one and evaluate it on a hold-out test set $\mathcal{D}_{\text{test}}\subset\mathcal{D}$. During all experiments, we train one probe per position (81 total, 12 seeds = 972 probes) and average the scores computed on a hold-out test over all. In the main text, we experiment with probing all cards (no matter if they have been observed). In Figure 11 and Figure 12 we show the opened vs. unopenend cards distinction.
@@ Appendix F Human Performance in Yōkai
We collected human game performance in a setting that adheres to the same setting that artificial agents face, i.e. participants played with 9 cards on a 9x9 grid in a two-player setting. Humans played in person. A pair of subjects was instructed on the rules verbally. Each pair of subjects was given a single training round to make sure that all participants were familiar with the rules. To make results as comparable as possible, humans were instructed to avoid any form of non-verbal communication. Each pair played five evaluation rounds, and we report mean return across games. Early termination was allowed according to the same rules as in the agent environment. The study was conducted under an umbrella agreement of the relevant institutional review board. Compensation was also handled under the rules of the relevant institution.
@@ Appendix G Example Games
Figure 13: We show an example game in which an agent wins in self-play. For each turn, we display the ground truth information at the top and the current observable information – essentially the agents’ observation – at the bottom. We also show how many card clusters exist at this point per turn. Hint cards are highlighted by a light-blue border. In turn 7, the agent finishes the final cluster of cards by correctly inferring that the card under the rbrb-hint must be red and not blue. In turn 8, agent 1 should have finished early for more reward, but did not recognise that fact and instead used the last hint card, thus finishing the game. Note that the agents tend to play quite conservatively. In turns 1 and 3, for instance, it gains knowledge of two separate green cards but only groups them in turn 5 after agent 1 places a corresponding gbgb-hint in turn 4.
@@
Figure 14: We show an example game in which an agent loses in self-play. For each turn, we display the ground truth information at the top and the current observable information – essentially the agents’ observation – at the bottom. We also show how many card clusters exist at this point per turn. Hint cards are highlighted by a light-blue border. Note that agent 0 in turn 1 finds a blue group early and manages to form a cluster quickly. Agent 0 confirms this by turn 3. In turn 6, the game is still winnable. However, because both agents over-focused on establishing a common ground around the blue cluster, they fail to realise this and in turn 7, agent 0 then blunders and misplaces the green card. From there, the game is lost under all circumstances. Additionally, note that in turns 4 and 6, agent 1 establishes the identity of all red cards but fails to recall that in turn 6, therefore also misplaying.
@@
We show an example game an agent wins in self-play in Figure 13 and one where it loses in Figure 14.
Appendix H 2 Player Cross-Play Matrices
Figure 15: Cross-play scores for 2 player self-play as featured in Table 2 ($\alpha=0.05$).
Figure 16: Cross-play scores for 2 player HE as featured in Table 2 ($\alpha=0.07$, $\lambda_{\text{GAE}}=0.85$).
Figure 17: Cross-play scores for 2 player OP as featured in Table 2.
Figure 18: Cross-play scores for 2 player OP ($\Phi_{\text{rot}}$ only) as featured in Table 2.
Figure 19: Cross-play scores for 2 player OP ($\Phi_{\text{colour}}$ only) as featured in Table 2.
Figure 20: Cross-play scores for 2 player OBL level 1 as featured in Table 2.
Figure 21: Cross-play scores for 2 player OBL level 2 as featured in Table 2.
Figure 22: Cross-play scores for 2 player OBL level 3 as featured in Table 2.
Figure 23: Cross-play scores for 2 player OBL level 4 as featured in Table 2.
Figure 24: Cross-play scores for 2 player OBL level 5 as featured in Table 2.
We show cross-play matrices for self-play (Figure 15, HE (Figure 16), Other-Play (for $\Phi$ in Figure 17, for $\Phi_{\text{rot}}$ in Figure 18 and for $\Phi_{\text{colour}}$ in Figure 19) and OBL (level 1 Figure 20, level 2 Figure 21, level 3 Figure 22, level 4 Figure 23 and level 5 Figure 24).
Appendix I Training Curves
Figure 25: SP training return for all two player OP results as shown in Table 3.
Figure 26: SP training curves for for two player OBL as shown in Table 2.
Figure 27: Training curves for all belief model levels for two player OBL as shown in Table 2.
We show relevant training curves in Figure 25, Figure 26 and Figure 27. All curves show convergence. As discussed in the main text, OP with $\alpha=0.08$ never learns any meaningful cooperation abilities and therefore shows a flat learning curve.
More insightful are the OBL learning curves. We can see that the belief model improves hugely from level 0 to level 1 which is expected given that the belief model now is trained on competent self-play behaviour. After this however, we surprisingly find that future belief models all level out at a final cross-entropy loss of about $0.36$. Consequently, higher levels of OBL policies also start to saturate at similar final return levels in self-play.